Circular No. 5876

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN OPHIUCHUS M. Della Valle, European Southern Observatory; and I. F. Mirabel and

B. Cordier, Service d’Astrophysique, Centre d’Etudes de Saclay, report: “We suggest as the probable optical counterpart of the x-ray nova GRS 1716–249 = GRO J1719–24 (cf. IAUC 5874) a star of B ∼ 17.1 that has been found, near the edge of the SIGMA/GRANAT error-box, on a CCD frame taken on Oct. 5.02 UT by J. Bonibaker (Leiden Observatory) with the Dutch 0.9-m telescope at La Silla. Inspection of the POSS B plate shows this star at B ∼ 21-21.5, implying a brightening by 4-4.5 magnitudes. A spectrum (range 420-720 mm, resolution ∼ 0.6 nm, low signal to noise) obtained on Oct. 5.05 UT by G. Stirpe (Osservatorio di Bologna) with the 1.5-m ESO telescope + B & C shows a somewhat featureless red continuum (type late G-early K) with narrow Hα that could be separated into two components, measured at 655.8 nm and 657.6 nm.”

I. F. Mirabel, Saclay; L. F. Rodriguez, Universidad Autónoma de México; and B. Cordier, Saclay, report: “Using the VLA we found on Oct. 3 near the edge of the SIGMA/GRANAT error circle a compact radio source at α = 17^{h}16^{m}32^{s}.52, δ = −24^{o}58′01′′.1 (equinox 1950.0; uncertainty 0′′.5). This position is coincident with that of the optical candidate reported above. The radio source has a flat spectrum with fluxes of 4.0 mJy at 6 cm and 4.3 mJy at 20 cm.”

NOTICE TO CONTRIBUTORS AND SUBSCRIBERS Further to IAUC 5874, we regret that, on receipt of line-charge pay-

ments, we shall be able to supply contributors in the future with only five copies, not twelve, of the printed Circulars containing their contributions.

Further to IAUC 5866, we point out that special (non-invoiced) subscribers who receive these Circulars only by e-mail will not easily be able to tell when their subscriptions will expire. This is one reason why electronic subscriptions were previously tied to subscriptions for the printed material. E-mail subscribers can also log in (at no extra charge) to our Computer Service, however, and they can thereby establish the expiry date for their subscriptions. The Computer Service is accessible by SPAN, telnet or modem. E-mail subscribers who have not received instructions on the use of the Computer Service should contact us (addresses above) for this information.

1993 October 6 Brian G. Marsden